RC meeting and Run Plan, Monday November 14th
From SBS wiki
Return to the RC Meeting Calendar
- 1 RC Meeting Information
- 2 Useful Links
- 3 Current Status
- 4 Run Plan
- 5 Subsystems Updates
- 6 Instructions for Shift Crew
- 7 Miscellaneous
RC Meeting Information
Time:at 3:30pm today
- Location: Counting house conference room
- Join ZoomGov Meeting
- Meeting ID: 161 903 2184
- Passcode: 260715
- Dial by your location
- +1 669 254 5252 US (San Jose)
- +1 646 828 7666 US (New York)
- +1 551 285 1373 US
- +1 669 216 1590 US (San Jose)
- 833 568 8864 US Toll-free
- Find your local number: https://jlab-org.zoomgov.com/u/acj3pTicMx
Accelerator and Hall Status
- We will take the Hall to "Controlled Access" tomorrow as soon as possible to accommodate all the work requests.
- Since Hall B is going to be down for a while we have requested MCC to change Wien angle tomorrow to maximize the polarization to Hall A. They have agreed to do so.
- We will stick to our previous plan to do a Moller measurement on Thursday. Should we take a mini spin dance measurement as well?
- If we stick to 45uA then MCC will not have to do a BLA calibration to Hall A BCM on Wednesday. They will have to do it if they need to send 60uA to both Hall A and C. Should we cancel our plan to do BCM calibration on Wednesday then?
- After a long downtime (4hrs+9hrs) we got beam back around 5pm yesterday. Then doing beam checkout and troubleshooting some BBGEM HV issue took another 3 hours. So, we could start production only around 8pm.
- Things have been smooth since then. MCC seems to be able to send us 45uA now.
- We had some DAQ related issues this morning but they are fixed now. Thanks to GEM folks and Alex.
Current Activity and Issues
- DAQ gets stuck somewhat frequently without obvious warning/error message. Shift crew gets confused and it may end up taking hours to fix. Is it possible to force CODA to throw better error messages? - Alex
- Target motion control freezes (either needs a iocha26 reboot, VMX controller reboot or a manual reset which requires an escorted access) which takes 2-3 hours to fix. Shift instructions updated to address this problem after investigation by Ellen - Arun
- Laser interlock reset box will be installed (in between the laser room and the entrance door) by Mark Taylor tomorrow. This is to avoid depolarization of the target after a laser trip and allow a more efficient way to reset the lasers without exposing the TO to lasers while not requiring them to be laser trained. The target-on-call will be notified each time and the lasers will be enabled by the TO after consultation. A key for the entrance door will be placed in the counting room key box for TOs - Arun
- Dump right Ion chamber unresponsive. For now director of operations agreed to allow running with just one dump IC. - J. Kowal and team will work on it (& Compton IC) tomorrow
- Continue taking production data
- 43uA of beam current | Use ps3=ps5=0, ps1=9, ps2=15, rest =-1 | One hour long runs
- Move the target to the pickup coil position and do an NMR sweep every 2/3 hours
- Return to step 1.
- Optics runs with sieve in - Should we put sieve in tomorrow before taking beam?
- Calibration runs on H2 - Do we have enough for this pass?
- Moller measurement - Thursday day shift (Mini Spin dance?)
- GEM luminosity study
- Polarization: 42.4% (12:17pm)
Quick Summary on the performance of every sub-system during GEn
Instructions for Shift Crew
- Keep an eye out for drifting BPM positions and TS1-BBHI rate. Post a screenshot of the relevant strip-charts and GUI twice every shift.
- If beam has been down due to accelerator issues perform a checkout of the beam profile by doing harp scans and check beam centering on target using the Carbon hole target, see Simona's procedure: procedure, If in doubt about whether you should go through the beam checkout procedure, call this number: 7578693131.
- Cross check the rate with what is expected: He3, H2
- Cross check with the expected ion chamber readings: log
- If holding field is changed for any reason, do NOT perform an NMR measurement.
Be sure to follow these procedures:
- Power down SBS, BB, and Detector HV before controlled access and allowing RadCon in the hall
- Communicate to the TO that the SBS and BB magnets will be powered down, so they can update the holding field
- See the beam recovery procedures at: HOW_TOs#Beam_Recovery_Procedure
- See the instructions to set the raster at: HOW_TOs#How_to_ask_MCC_to_set_the_desired_raster_size_of_6_mm
Experimental progress (for MCC planning meeting)
- 7.8% of accumulated charge for production data (18.18 Coulombs out of 232)
- 35% Experimental run time (50 out of 141 days)
- Long term "to do" items; not urgent, some can be done when 3He target not available
- Luminosity scans with GEMs. See procedure.
- Opportunistic BCM calibration (higher current) before pass change?
- INFN LV Scan and DAQ Stability - maybe 20-30 minutes if nothing else can run
- Work and testing on LV supplies was performed during opportunistic access. See log.
- Misc. GEM commissioning items: https://sbs.jlab.org/wiki/index.php/CommissioningPlan#BB.2FSBS_GEM_Commissioning_.28Continuation.29